HOW DO YOU FIND THE ALTITUDE OF A TRIANGLE THAT HAS A BASE OF 7 INCHES AND A SURFACE AREA OF 21 SGUARE INCHES?
Answer by ankor@dixie-net.com(22740)

HOW DO YOU FIND THE ALTITUDE OF A TRIANGLE THAT HAS A BASE OF 7 INCHES AND A SURFACE AREA OF 21 SGUARE INCHES?
:
You know the area of a triangle: .5*b*h = A
:
Substitute for b and A:
.5*7*h = 21
3.5h = 21
h = 21/3.5
h = 6 inches is the altitude
:
Check the area using these values
A = .5 * 7 * 6
A = 21 sq inches